# Topology of Discrete Quantum Feedback Control

### v1

@@ III.1 Setup
$\Lambda:=\{\vec{R}_{\vec{i}}=\sum_{\lambda=1}^{d}i_{\lambda}\vec{a}_{\lambda};\ \vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1\}$ (18)
be a set of lattice sites, where $\vec{a}_{\lambda}\ (\lambda=1,\cdots,d)$ denote the primitive lattice vectors. Each lattice site $\vec{R}_{\vec{i}}$ is specified by a set of integers $\vec{i}=(i_{1},\cdots,i_{d})$. The total number of unit cells is given by $N_{\mathrm{cell}}:=L_{1}L_{2}\cdots L_{d}$. The Hilbert space of the system is spanned by a basis {|$\{\ket{\vec{i},a}\}$ ($\vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1$, and $a=1,\cdots,D_{\mathrm{loc}}$), where |$\ket{\vec{i},a}$ denotes the quantum state at site $\vec{i}$, and the index $a$ labels additional degrees of freedom such as spin, orbit, sublattice, etc. The dimension of the Hilbert space in each unit cell is denoted by $D_{\mathrm{loc}}$, and thus the dimension of the entire Hilbert space $\mathcal{H}$ of the system is given by $D=N_{\mathrm{cell}}D_{\mathrm{loc}}$. (Note that we consider a one-body problem here.) The density matrix of the system can be expanded in terms of the orthonormal basis as
@@
which we shall refer to as the matrix representation of $\mathcal{E}$ (see Appendix A.1). Throughout this paper, we use a symbol with a tilde to indicate an operator acting on the doubled Hilbert space.
@@
for arbitrary $\hat{\rho}$, where $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ are translation operators:
@@
Exploiting the translational symmetry, we obtain the momentum-space representation of $\tilde{\mathcal{E}}$ as (see Appendix A.2 for derivation)
@@ III.2 Topology of quantum channels with local Kraus operators
We are now in a position to discuss topology of feedback control. We consider a homotopy between CPTP maps, that is, two CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ are homotopic to each other if and only if there exists a continuous one-parameter family $\Xi_{s}$ of CPTP maps (called a homotopy [81]) that preserves the spectral gap and satisfies $\Xi_{0}=\mathcal{E}_{0}$ and $\Xi_{1}=\mathcal{E}_{1}$. Intuitively speaking, the existence of a homotopy between two CPTP maps means that they can continuously be deformed to each other. Here, the spectral gap can be a point gap or a line gap according to the general discussion of topology of non-Hermitian matrices [33, 34]. Specifically, a point gap at $\xi=\xi_{\mathrm{PG}}$ means that a CPTP map under consideration does not have an eigenvalue $\xi_{\mathrm{PG}}$. Similarly, a line gap for a line $l_{\mathrm{LG}}$ means that a CPTP map does not have an eigenvalue on $l_{\mathrm{LG}}$. Here we focus on the case of a point gap.
@@
Here we note that the Bloch matrix is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. Since the size of the matrix diverges in the limit of $N_{\mathrm{cell}}\to\infty$, the definition of a topological invariant of a Bloch matrix is nontrivial. Here we show that this problem can be circumvented by imposing locality on the Kraus operators. We assume that the Kraus operators are local in space and have a finite support. Physically, this implies that the interaction between the system and the measurement apparatus (the feedback controller) is local. Let $\ell$ be the diameter of the support of the Kraus operators, and let $R$ be the range of the Kraus operators. Namely, we assume
@@
which indicates that the Bloch matrix $X(\vec{k})$ can be truncated without changing its nonzero eigenvalues and replaced by a matrix $X_{\mathrm{trunc}}(\vec{k})$ with a finite dimension that does not increase with increasing the system size (see Fig. 3). Hence, topology of CPTP maps can be diagnosed by that of the truncated Bloch matrix $X_{\mathrm{trunc}}(\vec{k})$, which is a finite-size matrix in the $N_{\mathrm{cell}}\to\infty$ limit. Topological feedback control is thus characterized by a topological invariant for the truncated Bloch matrix.
@@
which shows that |$\ket{\vec{i},a}\bra{\vec{j},b}$ | with $|\vec{i}-\vec{j}|>\ell$ is an eigenoperator of the CPTP map with eigenvalue zero. Thus, the truncation of a Bloch matrix and the highly degenerate zero eigenvalues can be understood as a consequence of local measurement, which erases quantum coherence between distant sites.
The existence of zero eigenvalues of quantum channels implies a crucial difference between dynamics described by those quantum channels and continuous time evolution described by the Lindblad quantum master equation. In fact, a quantum channel $\mathcal{E}$ that has a zero eigenvalue cannot be expressed in terms of any generator $\mathcal{L}$, i.e.,
@@ III.3 Stable equivalence of quantum channels
where $\hat{P}_{\vec{j},a}:=\ket{\vec{j},a}\bra{\vec{j},a}$ | is the projection operator. The projective measurement channel (40) plays a role of “trivial bands” in the sense that all of its nonzero eigenvalues are unity (see Appendix C). We note that the projection operators $\hat{P}_{\vec{j},a}$ are local in space and therefore in accordance with the discussion in Sec. III.2. We define the direct sum of a quantum channel $\mathcal{E}(\hat{\rho})=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}$ for the system and the projective measurement channel (40) for the extra bands by
@@
Figure 4: Stable equivalence of CPTP maps. A CPTP map with the eigenspectrum shown by blue areas in the top-left figure is not homotopic to the other CPTP map with the eigenspectrum shown by blue and green areas in the bottom figure, since the numbers of bands (eigenvalues) are different. However, they can be made homotopic if we add a trivial CPTP map with the eigenspectrum shown in the top-right figure.
@@ IV.1 Chiral Maxwell’s demon
Having established the general framework of topology of quantum channels, here we present a prototypical example of topological quantum feedback control characterized by a topological invariant of the Bloch matrix of a quantum channel. Our model is given by a quantum version of Maxwell’s demon in Ref. [54] (see also Ref. [82]). We consider a particle hopping on a one-dimensional (1D) lattice with $L$ sites, where a quantum state at site $i$ is denoted by |$\ket{i}$. We first perform a projective measurement of the position of this particle with measurement operators $\hat{M}_{m}=\ket{m}\bra{m}$ | $(m=1,\cdots,L)$. If the measurement outcome is $j$, we raise the potential at site $j-1$ and perform the feedback control with the Hamiltonian
@@
where $J\in\mathbb{R}$ is the hopping amplitude and $V>0$ is the height of the potential. The unitary operator for the feedback control is given by $\hat{U}_{j}=\exp[-i\hat{H}_{j}\tau]$, where $\tau$ denotes the duration of the feedback control. The potential barrier placed at site $j-1$ prevents the particle from going to the left, thereby achieving unidirectional (chiral) transport of the particle to the right (see Fig. 5). Here, we do not need any external field to induce the current in the system, but the information obtained from measurement is used to transport the particle. The feedback control thus rectifies quantum fluctuations of measurement outcomes and convert them into a current in a manner similar to an information ratchet [83]. We therefore refer to this model as the chiral Maxwell’s demon.
Figure 5: Schematic illustration of the chiral Maxwell’s demon. A quantum-mechanical particle hops on a one-dimensional lattice with hopping amplitude $J$. We perform a projective measurement of the position of a particle. If we find the particle at site $j$, we raise the potential at site $j-1$ to prevent the particle from hopping to the left.
@@
is the probability of finding a particle at site $j+r$ after feedback control with measurement outcome $j$. This probability does not depend on $j$ because of the translational invariance. The zero eigenvalues $\xi_{n\neq 0}(k)$ are due to the projective measurement in which quantum coherence between different sites vanishes: $\hat{M}_{m}\ket{i}\bra{j}\hat{M}_{m}^{\dagger}=0\ (i\neq j)$.
By using the eigenvalues, a topological invariant of the quantum channel is given by the winding number
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap [33]. Here, the eigenvalues $\xi_{n}(k)=0\ (n\neq 0)$ do not contribute to the winding number and thus are excluded from Eq. (46). The winding number takes an integer value and it is invariant under continuous deformation of a feedback control protocol as long as the point gap at $\xi_{\mathrm{PG}}$ is not closed. This is consistent with the $\mathbb{Z}$ classification of non-Hermitian matrices in 1D systems according to stable equivalence [33].
@@ IV.2 Eigenspectrum of a quantum channel under the periodic and open boundary conditions
In Fig. 6, we show the eigenspectrum of a quantum channel for the chiral Maxwell’s demon under the periodic boundary condition (PBC). Beside the highly degenerate zero eigenvalues $\xi_{n\neq 0}(k)$, the loop structure formed by the eigenvalue $\xi_{0}(k)$ can be seen. The winding number is $w(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. The eigenvalue $\xi_{0}(k)$ is determined from the probability distribution of the position after feedback control as shown in Eq. (43). In fact, the eigenvalue $\xi_{0}(k)$ is nothing but the characteristic function of the probability distribution $\{p(r)\}$. The chiral transport caused by the feedback control increases the probability of a particle being found at site $r>0$ and makes the probability distribution $\{p(r)\}$ asymmetric with respect to $r>0$ and $r<0$. Thus, the winding number of the quantum channel is intimately related to the chiral transport achieved by the feedback. This physical interpretation is in close analogy with the chiral edge state in the quantum Hall effect and the Hatano-Nelson model for non-Hermitian topological phases [33, 84, 85, 86, 87].
@@
Figure 6: Eigenspectrum of a quantum channel for the chiral Maxwell’s demon. A blue curve and a blue dot constitute the eigenspectrum under the periodic boundary condition in the limit of infinite system size, and orange dots show the eigenspectrum under the open boundary condition with system size $L=50$. The unit of energy is set to $J=1$, and we take $V=+\infty$ for simplicity. The duration of feedback is $\tau=1$.
@@
Nontrivial topology around a point gap under the PBC implies the emergence of a non-Hermitian skin effect under the open boundary condition (OBC) [33, 72, 73, 38, 74]. In Fig. 6, we show the eigenspectrum of the quantum channel under the OBC by orange dots. The eigenspectrum under the OBC is drastically different from that under the PBC, which is a salient feature of the non-Hermitian skin effect. The steady state of the quantum channel under the PBC is uniform in space, while it is localized near the right edge of the system under the OBC [see Fig. 7(a) and (b)]. In Fig. 7(c) and (d), we show a typical behavior of right and left eigenmodes in the OBC. The localization of the right and left eigenmodes at the opposite boundaries of the system indeed confirms the non-Hermitian skin effect, which is a consequence of the chiral transport induced by the chiral Maxwell’s demon.
@@
We note that the localization of right and left eigenmodes $\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}$ at opposite edges implies that the overlap $\langle\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}\rangle$ decreases exponentially as a function of the system size. From the eigemode expansion (14), this inditates that the expansion coeffecient $c_{n}$ becomes exponentially large, which significantly affects the relaxation dynamics of the system towards the steady state under repeated applications of a quantum channel [88, 89]. Specifically, if the non-Hermitian skin effect takes place, the relaxation time in the OBC generically diverges in the infinite-size limit.
@@
Figure 7: Magnitude $|\rho_{i,j}|$ of matrix elements of an eigenmode $\hat{\rho}_{n}=\sum_{i,j}\rho_{i,j}\ket{i}\bra{j}$ | plotted in the position basis for the system size of $L=20$ with the hopping amplitude $J=1$, the height of the potential $V=\infty$, and the duration $\tau=1$ of feedback control. Steady state of the chiral Maxwell’s demon under the PBC (a) and under the OBC (b). Typical right eigenmode (c) and left eigenmode (d) under the OBC. In (c) and (d), eigenmodes with eigenvalue 0.334 are shown.
IV.3 Topological phase transition

### v2

@@ III.1 Setup
$\Lambda:=\left\{\vec{R}_{\vec{i}}=\sum_{\lambda=1}^{d}i_{\lambda}\vec{a}_{\lambda};\ \vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1\right\}$ (18)
be a set of lattice sites, where $\vec{a}_{\lambda}\ (\lambda=1,\cdots,d)$ denote the primitive lattice vectors. Each lattice site $\vec{R}_{\vec{i}}$ is specified by a set of integers $\vec{i}=(i_{1},\cdots,i_{d})$. The total number of unit cells is given by $N_{\mathrm{cell}}:=L_{1}L_{2}\cdots L_{d}$. The Hilbert space of the system is spanned by an orthonormal basis set {|$\{\ket{\vec{i},a}\}$ ($\vec{i}=(i_{1},\cdots,i_{d}),i_{\lambda}=0,\cdots,L_{\lambda}-1$, and $a=1,\cdots,D_{\mathrm{loc}}$), where |$\ket{\vec{i},a}$ denotes a quantum state at site $\vec{i}$ with the index $a$ labeling additional degrees of freedom such as spin, orbit, sublattice, etc. The dimension of the Hilbert space in each unit cell is denoted by $D_{\mathrm{loc}}$, and thus the dimension of the entire Hilbert space $\mathcal{H}$ of the system is given by $D=N_{\mathrm{cell}}D_{\mathrm{loc}}$. (Note that we consider a one-body problem here.) The density matrix of the system can be expanded in terms of the orthonormal basis set as
@@
which we shall refer to as the matrix representation of $\mathcal{E}$ (see Appendix A.1). Throughout this paper, we use a symbol with a tilde to indicate an operator that acts on the doubled Hilbert space.
@@
for arbitrary $\hat{\rho}$, where $\hat{T}_{\lambda}\ (\lambda=1,\cdots,d)$ is a translation operator:
@@
We use the translational symmetry to obtain the momentum-space representation of $\tilde{\mathcal{E}}$ as (see Appendix A.2 for derivation)
@@ III.2 Topology of quantum channels with local Kraus operators
We are now in a position to discuss topology of feedback control. We consider a homotopy between CPTP maps, that is, two CPTP maps $\mathcal{E}_{0}$ and $\mathcal{E}_{1}$ are homotopic to each other if and only if there exists a continuous one-parameter family $\Xi_{s}$ of CPTP maps (called a homotopy [81]) that preserves the spectral gap and satisfies $\Xi_{0}=\mathcal{E}_{0}$ and $\Xi_{1}=\mathcal{E}_{1}$. Intuitively speaking, the existence of a homotopy between two CPTP maps means that they can continuously be deformed to each other. Here, the spectral gap is either a point gap or a line gap according to the general discussion of topology of non-Hermitian operators [33, 34]. Specifically, a point gap at $\xi=\xi_{\mathrm{PG}}$ means that a CPTP map under consideration does not have an eigenvalue $\xi_{\mathrm{PG}}$. Similarly, a line gap for a line $l_{\mathrm{LG}}$ means that a CPTP map does not have an eigenvalue on $l_{\mathrm{LG}}$. Here we mostly focus on the case of a point gap and discuss the case of a line gap in Sec. VII.5.
@@
Here we note that the Bloch matrix is an $N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}\times N_{\mathrm{cell}}D_{\mathrm{loc}}^{2}$ matrix. Since the size of the matrix diverges in the limit of $N_{\mathrm{cell}}\to\infty$, the definition of a topological invariant of a Bloch matrix is nontrivial. Here we show that this problem can be circumvented by imposing locality on the Kraus operators. We assume that the Kraus operators are local in space and have a finite support. Physically, this implies that the interaction between the system and the measurement apparatus (the feedback controller) is local. Let $\ell$ be the diameter of the support of the Kraus operators, and let $R$ be the range of locality of the Kraus operators. Namely, we assume
@@
which indicates that the Bloch matrix $X(\vec{k})$ can be truncated without changing its nonzero eigenvalues and replaced by a matrix $X_{\mathrm{trunc}}(\vec{k})$ with a finite dimension that does not increase with increasing the system size (see Fig. 3). Hence, topology of CPTP maps can be diagnosed by that of the truncated Bloch matrix $X_{\mathrm{trunc}}(\vec{k})$, which remains to be a finite-size matrix in the $N_{\mathrm{cell}}\to\infty$ limit. Topological feedback control is thus characterized by a topological invariant for the truncated Bloch matrix.
@@
which shows that |$\ket{\vec{i},a}\bra{\vec{j},b}$ | with $|\vec{i}-\vec{j}|>\ell$ is an eigenoperator of the CPTP map with eigenvalue zero. Thus, the truncation of a Bloch matrix and the highly degenerate zero eigenvalues can be understood as a consequence of local measurement, which projects out quantum coherence between distant sites.
The existence of zero eigenvalues of quantum channels implies a crucial difference between the dynamics described by those quantum channels and continuous time evolution described by the Lindblad quantum master equation. In fact, a quantum channel $\mathcal{E}$ that has a zero eigenvalue cannot be expressed in terms of any generator $\mathcal{L}$, i.e.,
@@
If Kraus operators are not strictly local but have an exponentially decaying tail, topology of a quantum channel may be discussed by approximating them with local ones because topology is insensitive to small deformations. Physically, this should hold true unless rare measurement outcomes with extremely small probabilities due to an exponentially decaying tail of wavefunctions significantly alter the physics. While this is expected to be true for most cases, it would be interesting to find the cases where rare events due to the exponentially small tails of Kraus operators play a significant role in feedback control. We do not further discuss those situations in this paper.
@@ III.3 Stable equivalence of quantum channels
where $\hat{P}_{\vec{j},a}:=\ket{\vec{j},a}\bra{\vec{j},a}$ | is the projection operator onto the $a$th internal degree of freedom at site $\vec{j}$. The projective measurement channel (40) plays a role of “trivial bands” in the sense that all of its nonzero eigenvalues are unity (see Appendix C). We note that the projection operators $\hat{P}_{\vec{j},a}$ are local in space in accordance with the discussion in Sec. III.2. We define the direct sum of a quantum channel $\mathcal{E}(\hat{\rho})=\sum_{m}\hat{K}_{m}\hat{\rho}\hat{K}_{m}^{\dagger}$ for the system and the projective measurement channel (40) for the extra bands by
@@
Figure 4: Stable equivalence of CPTP maps. A CPTP map with the eigenspectrum shown by blue areas in the top-left figure is not homotopic to the other CPTP map with the eigenspectrum shown by blue and green areas in the bottom figure, since the numbers of bands (eigenvalues) are different. However, they can be made homotopic if we add a trivial CPTP map with the eigenspectrum shown in the top-right figure so that the numbers of eigenvalues become equal to each other.
@@ IV.1 Chiral Maxwell’s demon
Having established the general framework of topology of quantum channels, here we present a prototypical example of topological quantum feedback control characterized by a topological invariant of the Bloch matrix of a quantum channel. Our model is a quantum counterpart of Maxwell’s demon in Ref. [54] (see also Ref. [82]). We consider a particle hopping on a one-dimensional (1D) lattice with $L$ sites, where a quantum state at site $i$ is denoted by |$\ket{i}$. We first perform a projective measurement of the position of this particle with measurement operators $\hat{M}_{m}=\ket{m}\bra{m}$ | $(m=1,\cdots,L)$. If the measurement outcome is $j$, we raise the potential at site $j-1$ and perform the feedback control with the Hamiltonian
@@
where $J\in\mathbb{R}$ is the hopping amplitude and $V>0$ is the height of the potential. The unitary operator for the feedback control is given by $\hat{U}_{j}=\exp[-i\hat{H}_{j}\tau]$, where $\tau$ is the duration of the feedback control. The potential barrier placed at site $j-1$ prevents the particle from going to the left, thereby achieving unidirectional (chiral) transport of the particle to the right (see Fig. 5). Here, we do not need any external field to induce the transport in the system; instead, the information gained from measurement is used to drive the particle unidirectionally. The feedback control thus rectifies quantum fluctuations in measurement outcomes and convert them into a unidirectional transport in a manner similar to an information ratchet [83]. We therefore refer to this model as the chiral Maxwell’s demon.
Figure 5: Schematic illustration of the chiral Maxwell’s demon. A quantum-mechanical particle hops on a one-dimensional lattice with hopping amplitude $J$. We perform a projective measurement of the position of the particle. If we find the particle at site $j$, we raise the potential at site $j-1$ to suppress the probability of the particle hopping to the left.
@@
is the probability of finding a particle at site $j+r$ after feedback control with measurement outcome $j$. This probability does not depend on $j$ because of the translational invariance. The zero eigenvalues $\xi_{n\neq 0}(k)$ are due to the projective measurement for which quantum coherence between different sites vanishes: $\hat{M}_{m}\ket{i}\bra{j}\hat{M}_{m}^{\dagger}=0\ (i\neq j)$.
By using the eigenvalues, the topological invariant of a quantum channel is given by the winding number
@@
where $\xi_{\mathrm{PG}}$ denotes the location of a point gap [33]. Here, the eigenvalues $\xi_{n}(k)=0\ (n\neq 0)$ do not contribute to the winding number (since $\partial_{k}\xi_{n}(k)=0$) and thus are excluded from Eq. (46). The winding number takes an integer value and it is invariant under continuous deformation of a feedback control protocol as long as the point gap at $\xi_{\mathrm{PG}}$ is not closed. This is consistent with the $\mathbb{Z}$ classification of non-Hermitian matrices in 1D systems according to stable equivalence [33].
@@ IV.2 Eigenspectrum of a quantum channel under the periodic and open boundary conditions
In Fig. 6, we show the eigenspectrum of a quantum channel for the chiral Maxwell’s demon under the periodic boundary condition (PBC). Beside the highly degenerate zero eigenvalues $\xi_{n\neq 0}(k)$, the loop structure formed by the eigenvalue $\xi_{0}(k)$ is evident. The winding number is $w(\xi_{\mathrm{PG}})=-1$ for a point gap $\xi_{\mathrm{PG}}$ inside the loop. The eigenvalue $\xi_{0}(k)$ is determined from the probability distribution of the position after feedback control as shown in Eq. (43). In fact, the eigenvalue $\xi_{0}(k)$ is nothing but the characteristic function of the probability distribution $\{p(r)\}$. The chiral transport caused by the feedback control increases the probability of a particle being found at site $r>0$ and makes the probability distribution $\{p(r)\}$ asymmetric with respect to $r>0$ and $r<0$. Thus, the winding number of the quantum channel is closely related to the chiral transport achieved by the feedback in analogy with the chiral edge state in the quantum Hall effect and the Hatano-Nelson model for non-Hermitian topological phases [33, 84, 85, 86, 87].
@@
Figure 6: Eigenspectrum of a quantum channel for the chiral Maxwell’s demon. A blue curve and a blue dot constitute the eigenspectrum under the periodic boundary condition for infinite system size, and orange dots show the eigenspectrum under the open boundary condition for system size $L=50$. Energy is measured in units of $J=1$, and we take $V=+\infty$ for simplicity. The duration of a feedback operation is $\tau=1$.
@@
Nontrivial topology around a point gap under the PBC implies the emergence of a non-Hermitian skin effect under the open boundary condition (OBC) [33, 72, 73, 88, 38, 74]. Orange dots in Fig. 6 show the eigenspectrum of the quantum channel under the OBC, which is drastically different from that under the PBC. Such an extraordinary sensitivity of the eigenspectrum to the boundary condition is a salient feature of the non-Hermitian skin effect. The steady state of the quantum channel under the PBC is uniform in space, while it is localized near the right edge of the system under the OBC [see Fig. 7(a) and (b)]. In Fig. 7(c) and (d), we show typical behavior of right and left eigenmodes under the OBC. The localization of the right and left eigenmodes at the opposite boundaries of the system indeed confirms the non-Hermitian skin effect, which is a consequence of the chiral transport induced by the chiral Maxwell’s demon.
As seen from Fig. 7(c) and (d), the localization length of a right eigenmode is, in general, different from that of the corresponding left eigenmode due to the difference between the dynamics near the right boundary and those near the left boundary. Specifically, the right eigenmode in Fig. 7(c) is less localized than the left eigenmode in Fig. 7(d) because of the reflection at the right boundary. This phenomenon arises from the fact that the quantum channel under the OBC is defined from the time evolution under the same boundary condition so that the complete positivity and trace preservation are ensured (see also Sec. VIII.2).
@@
We note that the localization of right and left eigenmodes $\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}$ at opposite edges implies that the overlap $\langle\hat{\rho}_{n}^{L},\hat{\rho}_{n}^{R}\rangle$ decreases exponentially as a function of the system size. From the eigemode expansion (14), this inditates that the expansion coeffecient $c_{n}$ becomes exponentially large, which significantly affects the relaxation dynamics of the system towards the steady state under repeated applications of a quantum channel [89, 90]. Specifically, if the non-Hermitian skin effect takes place, the relaxation time under the OBC generically diverges in the infinite-size limit [90].
@@
Figure 7: Magnitude $|\rho_{i,j}|$ of matrix elements of an eigenmode $\hat{\rho}_{n}=\sum_{i,j}\rho_{i,j}\ket{i}\bra{j}$ | plotted in the position basis for the system size of $L=20$ with the hopping amplitude $J=1$, the height of the potential $V=+\infty$, and the duration $\tau=1$ of feedback control. Steady states of the chiral Maxwell’s demon under the PBC and the OBC are shown in (a) and (b), respectively. Panels (c) and (d) show typical right and left eigenmodes under the OBC with eigenvalue 0.334 (see Fig. 6).
IV.3 Topological transition